OpenStudy (elusive):

Please... tan(sin^-1(-4/5))

OpenStudy (dumbcow):

Let \[\theta = \sin^{-1} (-\frac{4}{5})\] Then \[\ sin \theta = -\frac{4}{5}\]  \[\tan \theta =\pm \frac{4}{x}\] where \[x^2 = 5^2 - 4^2 = 3^2\]

OpenStudy (elusive):

what does the negative sign do? in -4/5

OpenStudy (elusive):

also, isn't that the unit circle? why is the hypothesis/radius 5? I thought it had to be 1

OpenStudy (shawn):

arcsine can only give you a value between -pi/2 and pi/2. So the angle is in the 4th quadrant

OpenStudy (shawn):

when you solve for x using Pythagiran Theorem, you should get x to be positive. I.e, x = +3 so tan(arcsin(-4/5)) = -4/3

OpenStudy (nerdrage29):

-4/3. Please help me on mine.

ganeshie8 (ganeshie8):

Hey

OpenStudy (elusive):

@ganeshie8 hey! omg I forgot to ping you

ganeshie8 (ganeshie8):

Let's try and do this in another way

ganeshie8 (ganeshie8):

Recall that tanx = sinx/cosx

OpenStudy (elusive):

What do you mean, do you still want my user?

ganeshie8 (ganeshie8):

Do you know how to express cosx in terms of sinx ?

OpenStudy (elusive):

how do you do that?

ganeshie8 (ganeshie8):

What trig identity do you remember the most ?

OpenStudy (elusive):

"Trig identity" comes up in the next unit, for me.

ganeshie8 (ganeshie8):

You must be knowing some

ganeshie8 (ganeshie8):

even if you're not supposed to know

OpenStudy (elusive):

1Introduction to Trigonometry 2Trigonometric Functions 3 Working with Trigonometric Functions 4 Trig Identities I'm on unit 3 right now.

OpenStudy (elusive):

Well. yeah. Are you talking about simple functions like sin/cos/tan/csc etc

OpenStudy (elusive):

I don't know how to 'express cosx in terms of sinx'

ganeshie8 (ganeshie8):

sin^2x + cos^2x = 1 have seen this identity before ?

OpenStudy (elusive):

Pythagorean!

ganeshie8 (ganeshie8):

Yes, if you're happy with that identity, you may use it to express cosx in terms of sinx

OpenStudy (elusive):

\[\cos(x) = \sqrt{1-\sin^2(x)}\]

ganeshie8 (ganeshie8):

Yep, So what's tanx in terms of sinx only ?

OpenStudy (elusive):

\[tanx =\frac{ \sin (x) }{ \sqrt{1-\sin^2(x)} }\]

OpenStudy (elusive):

No?

ganeshie8 (ganeshie8):

Yep

ganeshie8 (ganeshie8):

use that formula to simplify the given expression

ganeshie8 (ganeshie8):

Replace x by sin^-1(-4/5)

OpenStudy (elusive):

Are we allowed to do that? It's just "sin(x)", not "sin^-1(x)"...

ganeshie8 (ganeshie8):

You're allowed do anything you want as long as you do it to both sides and follow rules. So go-ahead

OpenStudy (elusive):

\[\tan(x) = \frac{ \sin(\sin^-1(-4/5)) }{ \sqrt{1-\sin^2(\sin^-1(-4/5))} }\]

OpenStudy (elusive):

? Seems more complicated..

ganeshie8 (ganeshie8):

You forgot to replace x in the lefylt hand side

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))) = \frac{ \sin(\sin^-1(-4/5)) }{ \sqrt{1-\sin^2(\sin^-1(-4/5))} }\]

ganeshie8 (ganeshie8):

It only looms complicated, but its really easy. sin^-1 eats sin

OpenStudy (elusive):

"eats"?

ganeshie8 (ganeshie8):

Yes, An inverse function undoes its function

ganeshie8 (ganeshie8):

Let me ask you a question

ganeshie8 (ganeshie8):

Suppose you took 10$ from your friend, To undo your debt, what must you do ?

OpenStudy (elusive):

Give back $10

ganeshie8 (ganeshie8):

Yes, so if take(x) is your function, then its inverse function is give(x)

ganeshie8 (ganeshie8):

what's the value of give(take(10)) ?

OpenStudy (agent0smith):

@ganeshie8 the value from whose POV? ;P @dumbcow's method works too, except sin^-1 only returns angles in quadrants 1 and 4, so the angle is only quadrant 4 (not in 3)

ganeshie8 (ganeshie8):

Yeah I should be more clear on that question : Suppose you had $100 in your pocket(x = 100). Then you took $10 : take(100) = 110 Then you gave back $10 : gave(110) = 100

ganeshie8 (ganeshie8):

The net effect of taking and then giving is nothing. It brings you back to your original state : gave(take(100)) = 100

OpenStudy (agent0smith):

That works better. The one you originally described give(take(10)) = 0 (for your own POV)

OpenStudy (elusive):

im sorry i had to leave yesterday.. wait.. im reading this now

OpenStudy (elusive):

okay.. so what happens next?

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))) = \frac{ \sin(\sin^-1(-4/5)) }{ \sqrt{1-\sin^2(\sin^-1(-4/5))} }\]

ganeshie8 (ganeshie8):

sin eats sin^-1 simply cancel them

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))) = \frac{ 1 }{ \sqrt{1-\sin} }\] ????

OpenStudy (elusive):

doesn't look right

OpenStudy (agent0smith):

The sin cancels out inverse sin. It does NOT cancel out the number inside the sin.

ganeshie8 (ganeshie8):

Also keeping mind, sin^2(x) is just another way to write (sin(x))^2

OpenStudy (agent0smith):

^Yeah don't forget that either \[\tan(\sin^-1(-4/5))) = \frac{ \sin(\sin^-1(-4/5)) }{ \sqrt{1-\left( \sin(\sin^-1(-4/5)) \right)^2} }\] I still think @dumbcow's method is the easier one, though

ganeshie8 (ganeshie8):

Yes, this is supposed to be an alternative

OpenStudy (elusive):

What's dumbcow's method?

OpenStudy (agent0smith):

The method we used in a question you did yesterday, @elusive... drawing a triangle.

ganeshie8 (ganeshie8):

I think you have almost finished the problem Cancel the sin, sin^-1 and see if you can simplify

OpenStudy (elusive):

The part I don't understand is.. Why are the triangles in quadrant 3 and 4? Because the sign is negative?

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))=\frac{ -4/5 }{ 1- -4/5^2}\]

ganeshie8 (ganeshie8):

Yes, sin is negative in 3, 4 quadrants. But this is only half of the answer

ganeshie8 (ganeshie8):

The other half had to do with domain of sin^-1

ganeshie8 (ganeshie8):

Looks you forgot the radical in the bottom

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))=\frac{ -4/5 }{ 41/25}\] how do you simplify further, now?

OpenStudy (agent0smith):

You forgot to square it... in the denominator, a negative number squared is positive

OpenStudy (elusive):

\[\tan(\sin^-1(-4/5))=\frac{ -4/5 }{ 9/25}\]

ganeshie8 (ganeshie8):

\[\dfrac{-4/5}{\sqrt{1-9/25}}\]

OpenStudy (elusive):

I would also like to know the triangle method.. I need to memorize some rules and properties?

OpenStudy (agent0smith):

@elusive the very first post in this question shows the triangle method. And you and I used it in another question of yours.

ganeshie8 (ganeshie8):

Not really, but you need to know some more about inverse functions first.

OpenStudy (elusive):

thank you, but how do you know which quadrant to put the triangles in??

OpenStudy (agent0smith):

You have to know/use the domains of the inverse functions to decide

ganeshie8 (ganeshie8):

Looking at sinx = -4/5, you figure out that x must be somewhere in 3rd or 4th quadrants

OpenStudy (elusive):

when graphing the inverse trig function, you just simply switch the x and y.. right? i think that's easy. I don't know what this thing we're doing right now, though.

ganeshie8 (ganeshie8):

That's the first step

OpenStudy (elusive):

The domains and range simply switch, right?

ganeshie8 (ganeshie8):

How do you know an inverse exists ?

OpenStudy (elusive):

humans create them

ganeshie8 (ganeshie8):

Heard of horizontal line test ?

OpenStudy (elusive):

yup

ganeshie8 (ganeshie8):

Yes, we create them and we want them to be well defined